### v1

# Regularized Adjusted Plus-Minus Models for Evaluating and Scouting Football (Soccer) Players using Possession Sequences

Abstract
This paper presents a novel framework for evaluating players in association football (soccer). Our method uses possession sequences, i.e. sequences of consecutive on-ball actions, for deriving estimates for player strengths. On the surface, the methodology is similar to classical adjusted plus-minus rating models using mainly regularized regression techniques. However, by analyzing possessions, our framework is able to distinguish on-ball and off-ball contributions of players to the game. From a methodological viewpoint, the framework explores four different penalization schemes, which exploit football-specific structures such as the grouping of players into position groups as well as into common strength groups. These four models lead to four ways to rate players by considering the respective estimate of each model corresponding to the player. The ratings are used to analyze the 2017/18 season of the Spanish La Liga. We compare similarities as well as particular use cases of each of the penalized models and provide guidance for practitioners when using the individual model specifications. Finally, we conclude our analysis by providing a domain-specific statistical evaluation framework, which highlights the potential of the penalized regression approaches for evaluating players.
1 Introduction
Evaluating player performances in professional football, sometimes referred to as soccer, has attracted a lot of interest in the last decade. Due to the low-scoring and complex nature of the sport, accurately measuring a player’s contribution to the game is a challenging task. However, evaluating individual performances in team sports is a fundamental task of increasing importance. While there is an external interest from fans (McHale et al., 2012), media (Kharrat et al., 2020), and the betting industry (Michels et al., 2023) another particularly important aspect nowadays is to use player ratings to inform scouting and retention decisions of teams while being compliant with budget constraints (Payyappalli and Zhuang, 2019; Pantuso and Hvattum, 2021).
In football, the decision-making process is traditionally realized by a qualitative assessment of players by professionals such as coaches, managers, or scouts. However, lately, there have been many attempts and advances toward a more objective approach to quantifying the impact of players on the game. For example, traditional statistics such as goals or pass completion percentage, which have been used extensively to analyze the strength of players, have been replaced by metrics such as expected goals (xG, cf. e.g. Robberechts and Davis, 2020; Anzer and Bauer, 2021) or passing models based on statistical approaches (Szczepański and McHale, 2016; Håland et al., 2019, 2020). In general, these novel data-driven methods can be categorized into two distinct groups: bottom-up and top-down rating systems (cf. Hvattum and Gelade, 2021). The former assesses player performances by assigning values to each action performed and then aggregating them for each player over the course of a relevant period (i.e. a match or a season). Top-down ratings on the other hand evaluate players by breaking down the whole team’s performance and distributing credit onto the players involved. Typically, bottom-up approaches use very granular data on an event level in order to assign values to either specific actions such as shots (xG) or passes or more generally any type of action (cf. Decroos et al., 2019). While bottom-up ratings are therefore able to assess a player’s quality down to each particular action performed by the athlete, they however lack the ability to account for team strengths as well as off-ball actions by teammates. In contrast, top-down ratings such as plus-minus (PM) models (see Hvattum, 2019 for an overview) rely only on information gained from specific segments of a match, which in general is much less detailed than event stream data. These approaches are by design not able to account for specific actions taken by a player. They are, however, able to incorporate effects of teammates.
In this work, we present a novel framework for evaluating player strength. On the surface, the methodology is similar to plus-minus (PM) models as we mainly use penalized regression models in order to infer a rating of players (cf. Kharrat et al., 2020; Hvattum, 2019). However, we extend the classical PM approach in two ways. First, we separate matches into possessions, which are sequences of consecutive on-ball actions. Thus, we consider much more granular segments of the game than the usual PM models, where a segment is defined as one where the same set of players is on the field. The segmentation in possessions leads to a data structure where it is possible to separate effects of being part of a possession, i.e. actually performing on-ball actions, from effects of only being on the pitch. From a domain-specific viewpoint, this is highly attractive as it is well known that off-ball actions have a substantial impact in soccer (cf. Spearman, 2018; Fernández et al., 2021). Using possessions data, it is also necessary to adjust the dependent variable that measures the outcome of a segment. In the classical PM approach the (normalized) goal difference in a segment is used. Since possessions are way shorter segments, goal difference is not appropriate as response variable. Instead, we focus on whether the possession ends in a goal or not as our outcome variable. Second, we provide a variety of penalization structures, which have not been considered in this context. Most of the classical plus-minus models use a ridge regression penalty, i.e. an $L_{2}$-type of penalty on the regression coefficient (cf. e.g. Kharrat et al., 2020). In a more innovative approach Pantuso and Hvattum (2021) adjust the penalty term by accounting for an age effect as well as a similarity effect within the shrinkage procedure. However, to the best of our knowledge, there has been no comparison or analysis of penalization structures beyond (adaptions of) the ridge penalty. We analyze several regularization structures that exploit the implicit grouping of players that arises from the different position groups on the field. In this way, our player evaluations are able to provide interesting and novel insights into player strengths where different penalties may be used for different use cases from practitioners. Finally, our last contribution aims at objectively evaluating ratings. In order to measure the relevancy of the novel modeling approaches presented in this work, we conduct a stastical evaluation specifically suited for football. The idea is similar to the validity experiments provided in Hvattum and Gelade (2021). That is, we measure whether a rating is superior to another rating by analyzing the ability of improving match results predictions using only the information from the rating. Besides comparing the ratings for the different regularized models, we also present comparisons to reference models using different kinds of data (ELO team ratings) and another modeling approach using a debiased machine learning approach (see Bajons, 2023).
The remainder of the paper is structured as follows. In Section 2 we describe the data used for this research and the necessary preprocessing steps. In Section 3 we provide details on the novel regularization structures and models used to determine player strength. Section 4 presents a comparison of the ratings as obtained from the different models as well as use cases and an approach to statistically evaluate the rating methods. Finally, we conclude the paper with a discussion in Section 5.
2 Data
We use raw event stream data from the Spanish La Liga season of 2017/18, openly available via figshare (Pappalardo et al., 2019), as a basis in order to create a dataset of possessions. A possession is defined as a sequence of consecutive on-ball events, which ends either with the opponent team gaining the possession or by an action of the referee, i.e. a foul, the ball leaving the pitch, or the end of a period. As the main goal is the valuation of players, it is necessary to extract information on which player was involved in a possession sequence. Extracting information on offensive players involved in a possession is straightforward from the data, however, information on the opposing players is desirable as well. Since event stream data only provides information on the on-ball actions, it is difficult to extract meaningful opponent player information. One way to incorporate such information is to simply record all players on the field for a possession. Thus the regression matrix $X_{mod}=(X_{N_{inv}},X_{N_{of}})$ in this work contains $N=N_{inv}+N_{of}$ columns and $N_{p}$ rows. Each of the involvement columns $N_{inv}$ is a binary variable for a specific player, that is 1 whenever the player was involved in the possession and 0 otherwise. Each of the $N_{of}$ columns indicates whether a specific player was on the field for the possession team (indicated by 1) or on the field for the defensive team (indicated by $-1$) [^2] In this work $N_{inv}=N_{of}$ corresponds to the number of players in the data set.. The number of rows $N_{p}$ depends on the number of possessions analyzed. As we are interested in deriving a player rating it makes sense to only consider valuable possessions, where we define a possession as valuable if it ends in the last third of the pitch (i.e. sufficiently close to the opponent’s goal). Note that because we are interested in evaluating players we do not extract possession-specific variables. From the data, it would be possible to extract spatial and temporal features such as the length and speed of a possession. Thus, if we were solely interested in predicting the success of a possession it would probably be important to take such variables into account. However, the underlying assumption is that such spatial or temporal effects are due to the players being involved in the possession as well as on the field. That is, while it may be the case that quick but long possessions increase the chance of a successful outcome it is due to the quality of players involved and on the field that it is even possible to execute such an attack.
Finally, a relevant response variable is extracted, which allows us to analyze the effect of players on possessions. Since goals are the most important part of the game of football, a goal indicator is taken as outcome variable. As football is a low-scoring game, one drawback is that this outcome variable is quite imbalanced. While there are other possibilities for outcome variables, such as ordinal type variables or hybrid approaches, it is beyond the scope of this paper to consider such cases.
The final data set used in this analysis covers possessions obtained from all of the 380 matches played in the 2017/18 season of the Spanish La Liga. Following our domain-specific intuition, only possessions that are comprised of more than 3 consecutive actions are considered. The only exceptions to this rule are possessions that start with a free kick or a corner kick, as they are relevant for scoring goals and often comprise only two actions (e.g. the corner kick and a shot). In most other circumstances, short possessions are a result of duels for the ball and are thus not necessarily directly relevant for detecting a player’s influence on scoring a goal from a possession. Discarding such possessions decreases the number to be analyzed from 141973 to $N_{p}=53233$ possessions. This, in particular, eliminates all the penalty kicks in our data [^3] It is common practice to separate goals scored from penalty kicks from goals scored within the flow of the game when analyzing the ability of football players, compare e.g. Robberechts and Davis, 2020 and Anzer and Bauer, 2021.. As we measure goals as outcome variable and these are sparse, a different definition of possessions or a different data truncation scheme may be considered in the future. For this work, we stick to the approach described, as this amounts to an increase in the percentage of possessions resulting in a goal from 0.7 % to 1.3 %. The final number of players in our data is 556 players, i.e. $N_{inv}=N_{of}=556$ and thus $N=1112$. On average the possessions contain $\approx 7.5$ consecutive actions with a median of $6$ consecutive actions. The average number of direct involvements for players in possession is 394.8. The distribution of the number of involvements is skewed to the right, which intuitively reflects the idea that only a few outstanding players have a lot of involvements whereas especially defenders are not as often involved in possessions. It is important to mention, that from a domain-specific viewpoint, several possible adjustments to the data setup could be considered. While we discuss some of these in more detail in Section 5, we emphasize that for the purpose of presenting the benefits of our novel extensions of top-down ratings, the setup is well-suited.
3 Methodology
With the possessions data at hand, the most naive and straightforward approach to deriving a player rating would be to use classic regression techniques and fit a generalized linear model to the data. Then the models’ regression coefficients could be used as a strength estimate of each player. However, as many studies point out (Gramacy et al., 2013; Sæbø and Hvattum, 2015; Hvattum, 2019; Kharrat et al., 2020), such an approach is not expected to work in the high-dimensional and sparse setup at hand, due to classical issues such as multicollinearity, overfitting and the high imbalance in outcome and independent variables.
In the following, we discuss several approaches to deriving a rating of football players from the processed event data that overcome the previously mentioned problems. The main class of models is regularized or penalized regression, often referred to as adjusted plus-minus models, which have been studied in the context of sports player ratings by many authors previously (e.g., Gramacy et al., 2013; Hvattum, 2019; Kharrat et al., 2020). We adapt the methodology to our data setup and discuss various forms of penalized regression models. For each of the models, first, a statistical description is provided and secondly, a detailed domain-specific interpretation of the penalization structure as well as practical implications and use cases for practitioners are discussed.
3.1 Ridge regression
One of the earliest approaches for regularization dating back to Hoerl and Kennard (1970) is ridge regression (sometimes also referred to as Tikhonov regularization). The ridge estimator is given
[…]
where $\ell(Y;X,\beta)$ represents the (negative) log-likelihood function for the binomial response variable $Y$ in the usual generalized linear model framework that we work in. In the case of a Gaussian likelihood function, an explicit solution for the above problem can be derived (see e.g. Hastie et al., 2013). For generalized linear models, the problem is typically formulated as a weighted least-squares problem which is then solved in an iterative fashion (cf. Friedman et al., 2010). The solution $\hat{\beta}_{ridge}$ of problem (1) is quite appealing as it shrinks the coefficients towards each other, where more shrinkage is put on the columns that explain less variance in the regression matrix $X$. That is, as we expect the response variable, i.e. goals scored from possession, to vary most in the directions of important players, this approach shrinks regular players more than strong (or weak) players. However, players with very similar strength will likely be shrunken towards a similar value. This is quite desirable from a perspective of a rating, as every player gets a rating that reflects his strengths in relation to every other player on the field.
Another particularly nice feature arises from the data setup. As the regression matrix $X=(X_{N_{inv}},X_{N_{of}})$ consists of involvement columns $N_{inv}$ of each player as well as of field indicator columns $N_{of}$, there are two sources of information on each player’s strengths. One is the direct information of how much a player $i$ contributes to specific possessions when involved, given by the corresponding (shrunken) estimate representing player $i$’s column in the sub-matrix $X_{N_{inv}}$. The other one is the indirect contribution to possessions, given by player $i$’s estimate from the sub-matrix $X_{N_{of}}$. From a practical perspective this is nicely interpretable: A player, even though potentially not involved in the possession, usually nevertheless has an influence on the game. A threatening forward for example may create space for other players, simply because the opposing team tries to cover the player. On the other hand, a strong goalkeeper who deflects a lot of shots also clearly influences offensive possessions of the opposing team. Such an effect may be measured from the estimate of indirect involvement of a player to a possession. The total strength of a player could then e.g. be measured by the sum of both contributions. In section 4 we discuss several methods of assembling the two strength estimates of a player in order to obtain an overall strength estimate of players.
3.2 Group lasso and exclusive lasso
A further very popular form of regularization is lasso regression, where instead of imposing an $L_{2}$-norm constraint on the coefficients $\beta$ (as is done in ridge regression), an $L_{1}$-norm penalty is used. The estimate is thus defined as
[…]
This penalty allows for implicit variable selection, as some of the estimates are not only shrunken towards zero but set exactly to zero (cf. e.g. Hastie et al., 2013). While such an effect is desirable in many situations, as variables with less influence are omitted from the regression model and only influential variables retain a value different from zero, in our setup it is suboptimal. Even though it may be possible to detect strong (or weak) players from the resulting coefficients, the majority of them will be set to zero and no interpretation is possible anymore. This is especially inconvenient for the case of recruiting players, as a team is usually looking for a tradeoff between a strong player and an affordable player. Furthermore, the appealing effect of interpreting direct and indirect involvement to possession as in the ridge regression case is not as easy to employ. In the following, we discuss variants of the lasso penalty, which are more useful for our player evaluation framework.
A particularly popular variant is the group lasso penalty (see e.g. Simon et al., 2013) given as
$\quad\min_{\beta}\ell(Y;X,\beta)+\lambda\sum_{l=1}^{m}\sqrt{p_{l}}\left\|\beta^{(l)}\right\|_{2},$ (3)
where the predictor variables are divided into $m$ groups. In equation (3), $\ell(Y;X,\beta)$ represents a scaled version of the negative log-likelihood and $p_{l}$ are weights representing the length of the $l$-th group specific estimate $\beta^{(l)}$ (cf. Simon et al., 2013). This penalty results in a shrinkage of the coefficients of complete groups to zero, i.e. on a group level it performs shrinkage similar to lasso. However, within each group, it does not perform variable selection, leading to a similar effect as ridge regression. Thus, if each group is of size one the group lasso results in the classical lasso. On the other hand, if there is only one group the penalty is similar to ridge regression.
In our application of rating football players, there is a natural grouping of players as usually, each player plays a specific position. While these positions may vary slightly depending on specific matches and tactics used, players can be categorized into at least 4 broad categories: Forward, midfielder, defender, and goalkeeper. The group lasso enables us to distinguish between players in each of these groups. On a group level, whole groups are shrunken simultaneously. If there were one irrelevant group (say for example the goalkeepers, whose main task is not to be involved in possession of high threat), then this group would be shrunken towards zero. Within a group however, the ridge-like penalty shrinks players of similar strength to a similar value, while less shrinkage is applied to strong (or weak) players. From an interpretational viewpoint, this is highly desirable. As the tasks on the field of each individual group vary, we do not expect comparisons between two groups, e.g. forwards and defenders, to be fair and instead might be more interested in within-group comparisons. Another dimension arises again from the distinction between direct and indirect involvement. Similar to the ridge regression case one can combine both aspects for the groups and perform within-group analyses.
Another variation of the lasso is the exclusive lasso (see e.g. Campbell and Allen, 2017), which solves the problem
$\quad\min_{\beta}\ell(Y;X,\beta)+\frac{\lambda}{2}\sum_{l=1}^{m}\left\|\beta^{(l)}\right\|_{1}^{2},$ (4)
where $\ell(Y;X,\beta)$ again represents a scaled version of an exponential family likelihood function and the regressors can be separated into $m$ different groups. The behavior of this penalty is analyzed in detail in Campbell and Allen (2017). As opposed to the group lasso, it can be obtained from equation (4) that between groups an $L_{2}$ (i.e. ridge) penalty is applied, while within each group an $L_{1}$ (i.e. lasso) penalty is applied. That is, on a group level the penalty does not imply shrinkage of complete groups, i.e. in each group at least one variable will be non-zero. Within groups however only the most relevant variables will be selected, i.e. the penalty tries to set most of the group members to zero as is usual for the lasso penalty. In the extreme case, exactly one non-zero variable of each group will be selected.
The application-specific use case of the penalty reveals more potential in scouting rather than in rating players. While one could again categorize the players into the above 4 categories based on their broad position on the pitch and analyze which players are particularly strong (resp. weak) in their respective group, the exclusive lasso motivates more interesting selection procedures. A particular use case would be a team that is interested in signing players who play a specific position and who are at a certain age level or below a certain market value. In this case, an individual group could be formed and the exclusive lasso would result in the selection of a few strong (resp. weak) players that fall into the prespecified type, while at the same time being able to account for the usual groups as defined by e.g. player position. More details on using the exclusive lasso for scouting are found in Appendix B for this work.
3.3 Generalized lasso
In the previous section, we discussed penalizing structures on prespecified groups and their usage in analyzing football players. However, a different approach is to set up the penalizing structure such that implicit grouping is achieved. In the context of analyzing sports teams, Masarotto and Varin (2012) proposed the ranking lasso, given by the problem
$\min_{\beta}\ell(Y;X,\beta)+\lambda\sum_{i<j}^{N}w_{ij}|\beta_{i}-\beta_{j}|.$ (5)
The penalty term here is a variation of the lasso that allows for the grouping of the estimated abilities of players into several equivalence classes. More generally the problem can be written as
$\min_{\beta}\ell(Y;X,\beta)+\lambda||D\beta||_{1},$ (6)
where $D$ is a matrix that allows for individual specification of lasso-type penalties. The penalization structure in equation (6) is termed generalized lasso (cf. e.g. Tibshirani and Taylor, 2011) and includes the ranking lasso as well as the more well-known fused lasso (Tibshirani et al., 2004).
Applying the generalized lasso penalty in the player evaluation setting is not completely straightforward. The aim is to cluster similar player contributions together, however as the structure of the regression matrix is $X=(X_{N_{inv}},X_{N_{of}})$, it does not make sense to consider the simple ranking lasso penalty. Instead, a more natural approach is to consider direct involvements separated from indirect ones. We first examine the structure of the matrix $D_{N_{inv}}$ for the coefficients corresponding to one of the sub-matrices of $X$, namely $X_{N_{inv}}$. The goal is to employ a penalty such that similar strength estimates are shrunken together, i.e. to apply the ranking lasso to the estimates of these columns. Thus the matrix $D_{N_{inv}}$ can be expressed as a block matrix
$D_{N_{inv}}=(A,B,C)=\left(\begin{array}[]{ccc}A_{1}&B_{1}&C_{1}\\ \vdots&\vdots&\vdots\\ A_{N-1}&B_{N-1}&C_{N-1}\end{array}\right).$ (10)
containing three blocks $A^{\top}=(A_{1}^{\top},\dots,A_{N-1}^{\top})$, $B^{\top}=(B_{1}^{\top},\dots,B_{N-1}^{\top})$ and $C^{\top}=(C_{1}^{\top},\dots,C_{N-1}^{\top})$. Each $A_{i}$ is a matrix of zeroes of dimension $(N_{inv}-i)\times(i-1)$, each $B_{i}$ is a matrix (actually a vector) of ones of dimension $(N_{inv}-i)\times 1$ and each $C_{i}$ the negative identity matrix of dimension $N_{inv}-i$. The same penalization structure can be applied to the coefficients of indirect involvement, i.e. these are also shrunken according to the ranking lasso penalty and thus $D_{N_{inv}}=D_{N_{of}}$. In order to obtain the full penalization matrix $D$, the two sub-matrices $D_{N_{inv}}$ and $D_{N_{of}}$ need to be concatenated suitably:
$D=\left(\begin{array}[]{cc}D_{N_{inv}}&\bm{0}\\ \bm{0}&D_{N_{of}}\end{array}\right).$ (13)
Indeed one can think of different approaches for the penalization matrix $D$, for example, if $D_{N_{of}}$ is the identity matrix, this would correspond to a classical lasso penalty on the indirect involvement coefficients. For the application to player evaluation, it makes however sense to stick to a structure where the coefficients of direct involvements in possessions are separated from the ones of indirect.
4 Computational results and comparison of approaches
4.1 Implementation
All of the models discussed in Section 3 are applied to the event data set described in Section 2. All computations are performed using the R statistical programming language (R Core Team, 2022) and we make use of several contributed packages. For the ridge regression, the glmnet function from the correspondent package (Friedman et al. (2010) and Tay et al., 2023) is used, the group lasso is fitted using the package SGL (Simon et al., 2019) and for the exclusive lasso the ExclusiveLasso package (Weylandt et al., 2018) is provided. In all cases, we use a suitable cross-validation procedure in order to select the optimal shrinkage parameter $\lambda$. For the generalized lasso, there is no R package readily available. Although there exists the genlasso package (Arnold and Tibshirani, 2022), this only provides an implementation for the case of Gaussian outcome variables.
We briefly describe our implementation for the binomial logistic regression. We use modern conic programming principles to solve the problem. In particular, equation (6) can be written as
$\min_{\bm{\beta}}-\big{(}\sum_{i=1}^{n}y_{i}\log(h_{\bm{\beta}}(x_{i}))+(1-y_{i})\log(1-h_{\bm{\beta}}(x_{i}))\big{)}+\lambda\sum_{j=1}^{m}w_{j}|D_{j}|,$ (14)
with logistic function $h_{\beta}(x)=1/\big{(}1+\exp(-\beta^{\top}x)\big{)}$ and pair-specific weights $w_{ij}$. In this case, $D_{j}$ is the $j$-th component of $D\bm{\beta}$, with $D\in\mathbb{R}^{m\times p}$. This problem can be reformulated as a conic program by expressing it in its equivalent epigraph form (see Boyd and Vandenberghe, 2004). Specifically, (14) can be written as
$\min_{(\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2})}$ $\sum_{i=1}^{n}t_{i}+\lambda r$ (15)
s.t. $\left(z_{i1},1,u_{i}-t_{i}\right)\in K_{\exp},$
$\left(z_{i2},1,-t_{i}\right)\in K_{\exp},$
$z_{i1}+z_{i2}\leq 1,$ $i=1,\dots,n,$
$-s_{j}\leq D_{j}\leq s_{j},$ $j=1,\dots,m,$
$r-w_{1}D_{1}-\dots-w_{m}D_{m}\geq 0.$
In the above formulation, the variables $\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2}$, and $\bm{u}$ are auxiliary variables. The initial problem is thus rewritten as a conic programming problem on an augmented set of variables $\Theta=(\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2})$. Such a problem can be solved using a convex optimization solver of choice. Note that in order to solve (15), only two types of conic cones are necessary, namely the linear cone and the exponential cone. We use the R-package ROI (Theußl et al., 2020) which allows to nicely handle such types of optimization problems in a flexible manner.
4.2 Results
4.2.1 Comparing ratings
In this section, the resulting ratings from the four approaches are analyzed and compared. We focus on the 2017/18 season of the Spanish La Liga and start by presenting the individual results and their implications and end with a comparison of the different models. In total, the methodology is applied to 556 players, and for each of the penalized regression approaches the estimates for direct involvement, indirect involvement as well as the sum thereof are analyzed.
Table 1: Top 5 of each position group (forward, midfielder, defender, goalkeeper) with regards to the sum of direct (Inv) and indirect involvement (OF) under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward L. Messi 0.04088 0.03344 0.07432 1 1 1
L. Suárez 0.03545 0.02547 0.06092 2 3 2
Cristiano Ronaldo 0.03228 0.01284 0.04511 3 44 3
Lucas Vázquez 0.01474 0.02482 0.03956 19 4 4
Iago Aspas 0.03169 0.00347 0.03516 5 157 7
Midfielder Paulinho 0.00942 0.02667 0.03609 40 2 5
I. Rakitić 0.01143 0.02391 0.03534 32 8 6
Pablo Fornals 0.01553 0.01568 0.03121 17 31 10
T. Kroos 0.01840 0.00647 0.02486 8 115 19
André Gomes 0.00218 0.02249 0.02467 136 10 20
Defender Nacho 0.00822 0.02456 0.03278 56 5 8
Piqué 0.00758 0.02445 0.03204 62 6 9
Marcelo 0.00707 0.02192 0.02899 66 11 13
Jordi Alba 0.00995 0.01794 0.02789 37 19 14
D. Godín 0.00289 0.02159 0.02448 122 14 21
Goalkeeper M. ter Stegen 0.00365 0.02344 0.0271 105 9 15
K. Navas 0.00607 0.01682 0.02289 73 25 24
J. Oblak -0.00055 0.02188 0.02133 305 12 32
Neto -0.00413 0.01647 0.01234 484 27 76
Iago Herrerín -0.00106 0.00858 0.00752 346 84 124
We first discuss the results from ridge regression. Table 1 shows the top 5 players for each of the four broad position groups discussed in Section 3 based on the sum of the coefficient for direct involvement (Inv) and indirect involvement (OF). While using the sum of the coefficients as a measure of player strength is to some degree arbitrary, the resulting table provides interesting insight into the behavior of the penalty. Considering the rank columns in Table 1 it becomes clear that the direct involvement coefficients favor offensive players such as forwards, while defenders and goalkeepers score better on the indirect involvements. This behavior is expected due to the way we set up our models. As the outcome variable measures whether a possession ends in a goal or not, clearly forwards should have a more direct impact on it since scoring goals is their main job. On the other hand, the main task of defenders and goalkeepers is to protect the team from receiving goals, which is rather measured through the indirect involvements. Thus, in order to measure the value of forwards and midfielders it may be more accurate to analyze their rank based on direct involvement as shown in Table 2. In the same manner, it may be more accurate to rank defenders and goalkeepers based on their indirect involvement coefficient as presented in Table 3.
Table 2: Top 10 forwards and midfielders with regards to direct involvement (Inv) in possessions under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward L. Messi 0.04088 0.03344 0.07432 1 1 1
L. Suárez 0.03545 0.02547 0.06092 2 3 2
Cristiano Ronaldo 0.03228 0.01284 0.04511 3 44 3
Maxi Gómez 0.03223 -0.01875 0.01348 4 544 68
Iago Aspas 0.03169 0.00347 0.03516 5 157 7
A. Griezmann 0.02288 0.00327 0.02615 6 163 16
Munir 0.01868 -0.01013 0.00856 7 493 110
C. Bacca 0.01833 -0.00198 0.01636 9 350 50
Rodrigo 0.01799 0.00070 0.01868 11 237 40
G. Bale 0.01668 0.00025 0.01692 13 255 46
Midfielder T. Kroos 0.01840 0.00647 0.02486 8 115 19
D. Wass 0.01815 -0.00181 0.01635 10 343 51
Samu Castillejo 0.01686 0.00495 0.02182 12 133 29
P. Sisto 0.01635 0.00335 0.01971 15 161 37
Pablo Fornals 0.01553 0.01568 0.03121 17 31 10
P. Hernández 0.01457 -0.01252 0.00205 20 512 211
L. Modrić 0.01392 0.00031 0.01423 23 251 66
Portu 0.01364 0.00063 0.01426 24 239 65
I. Rakitić 0.01143 0.02391 0.03534 32 8 6
Mikel Oyarzabal 0.01076 0.00094 0.0117 35 230 81
Table 3: Top 10 defenders and goalkeeper with regards to indirect involvement (OF) in possessions under the rating derived from ridge regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Defender Nacho 0.00822 0.02456 0.03278 56 5 8
Piqué 0.00758 0.02445 0.03204 62 6 9
J. Murillo -0.00177 0.02394 0.02217 389 7 25
Marcelo 0.00707 0.02192 0.02899 66 11 13
D. Godín 0.00289 0.02159 0.02448 122 14 21
Nélson Semedo 0.00325 0.01993 0.02318 113 15 23
J. Giménez 0.00217 0.01959 0.02176 138 16 30
Víctor Ruíz 0.00453 0.01950 0.02403 94 17 22
L. Hernández -0.00389 0.01855 0.01466 480 18 61
Jordi Alba 0.00995 0.01794 0.02789 37 19 14
Goalkeeper M. ter Stegen 0.00365 0.02344 0.0271 105 9 15
J. Oblak -0.00055 0.02188 0.02133 305 12 32
K. Navas 0.00607 0.01682 0.02289 73 25 24
Neto -0.00413 0.01647 0.01234 484 27 76
Fernando Pacheco -0.00645 0.01235 0.00590 530 52 143
Iago Herrerín -0.00106 0.00858 0.00752 346 84 124
Rubén Blanco -0.00247 0.00771 0.00523 432 95 155
Y. Bounou -0.00492 0.00653 0.00161 504 114 222
Kiko Casilla -0.00222 0.00567 0.00345 416 127 185
C. Pantilimon -0.00167 0.00467 0.00300 386 137 196
Next, the results from the group lasso regression are analyzed. For this exercise, the columns are split into 8 groups. The players’ direct and indirect involvements are grouped into the 4 player positions, resulting in 4 direct involvement groups and 4 indirect involvement groups. Table 4 is the analog of Table 1, i.e. it displays the top 5 players for each of the four positions under the group lasso penalty using the sum of the coefficients of direct (Inv) and indirect (OF) involvements.
Table 4: Top 5 of each position group (forward, midfielder, defender, goalkeeper) with regards to the sum of direct (Inv) and indirect involvement (OF) under the rating derived from group lasso regression.
Score Rank
Position Player Inv OF Sum Inv OF Sum
Forward Iago Aspas 0.74980 0.02308 0.77288 1 162 1
L. Messi 0.61590 0.15345 0.76934 3 9 2
Cristiano Ronaldo 0.68337 -0.01686 0.66650 2 389 3
L. Suárez 0.59205 0.02455 0.61660 4 155 4
Juanmi 0.49060 0.00323 0.49382 7 256 5
Midfielder Portu 0.40491 0.04977 0.45468 11 75 6
D. Wass 0.38755 -0.01856 0.36898 12 394 10
T. Inui 0.28829 0.03389 0.32218 18 119 15
Zurutuza 0.12923 0.13300 0.26223 65 11 23
Samu Castillejo 0.22036 0.0269 0.24725 30 141 27
Defender Nacho 0.05847 0.16819 0.22666 111 5 31
Víctor Ruíz 0.04315 0.16532 0.20847 131 7 35
Aritz Elustondo 0.11675 0.08441 0.20116 74 36 38
H. Moreno 0.08787 0.10561 0.19348 85 19 41
José Ángel 0.05719 0.10806 0.16525 113 17 50
Goalkeeper K. Navas 0.11321 0.0232 0.13642 76 159 72
J. Cillessen -0.01402 0.06396 0.04995 343 58 143
Toño 0.00000 0.04072 0.04072 226 103 157
Oier 0.00766 0.02601 0.03367 194 147 170
Jaume 0.01657 0.01000 0.02657 182 218 188
The general effect of the group lasso penalty is similar to the ridge regression, i.e. offensive players (forwards and midfielders) load higher on the coefficient of direct involvement and defensive players (defenders and goalkeepers) exhibit a tendency to load higher on the coefficient of indirect involvement. However, it seems that the effects are more pronounced, especially for the direct involvements. This is reasonable since the idea of the group lasso is to shrink whole groups and thus less relevant groups, such as the defensive players, are penalized more severely. Figure 1 analyses the relationship between the ridge estimates and the group lasso estimates in more detail. In terms of correlation between the two rating methods, it seems that there is the most agreement for the coefficients of direct involvements rather than the indirect involvements. Furthermore, the ratings of the offensive players are more highly correlated than the ones for the defensive players. Overall the coefficients seem to be similar in both types of involvement in possessions as in general, the correlation is greater than 0.75. The sum of the coefficients however exhibits the least correlation between the two ratings, suggesting that it may be worth studying different approaches of how to combine the two individual contributions to possession than simply summing them.
Figure 1: Correlation between the ratings from ridge regression and group lasso regression. Correlation of the scores (x-axis) and the ranks (y-axis) are displayed where each of the 4 position groups (forwards, midfielders, defenders, goalkeepers) are colored differently and each of the involvement types (Inv, OF, Sum) is displayed by a different symbol.
The results from the exclusive lasso are presented in Table 5 and Table 6. Again the same 8 groups as for the group lasso penalty are considered. As the exclusive lasso performs lasso-like shrinkage within the group but selects at least one relevant representative from each group, most of the players are shrunken to zero. However, the general pattern is similar to the two previous approaches. When analyzing direct involvement as in Table 5 offensive players score better than defensive ones. In fact, only one of the 3 selected defenders has a positive coefficient estimate and both estimates of the selected goalkeepers are negative. The coefficient estimates for the indirect involvement are displayed in Table 6. In this case, most of the defenders obtain positive values and the top player according to indirect involvement to possessions is a goalkeeper. Furthermore, the estimates of the selected midfielders are also positive, however, all of them are considered to primarily play in the center with strong defensive roles. Only two forwards are selected with only Lionel Messi obtaining a positive estimate of indirect involvement. Interestingly, Messi is also the only player selected with regards to both, involvement categories, and at the same time being the top selected player with regards to direct involvement. This suggests that his appearance on the field has a substantial impact on the game. As already mentioned, the results from the exclusive lasso are not optimal for rating or ranking players as only a few relevant players are not set exactly to zero. This precludes the possibility of comparing any set of players that may be of interest. In Appendix B, we discuss how the exclusive lasso penalty may be used for scouting players by defining groups of players that are of interest. In Section 4.3 we discuss a different approach for defining groups which is more useful for comparing the validity of the developed ratings.
Table 5: Selected players from each position group with regards to direct involvement (Inv) in possessions under the exclusive lasso penalty.
Score Rank
Position Player Inv OF Inv OF
Forward L. Messi $0.302$ $0.002$ $1$ $12$
L. Suárez $0.109$ $0.000$ $3$ $13$
Cristiano Ronaldo $0.043$ $0.000$ $6$ $13$
Maxi Gómez $0.033$ $0.000$ $7$ $13$
Midfielder T. Kroos $0.128$ $0.000$ $2$ $13$
Samu Castillejo $0.067$ $0.000$ $4$ $13$
D. Wass $0.064$ $0.000$ $5$ $13$
Defender Barragán $0.005$ $0.000$ $8$ $13$
Gabriel -$0.028$ $0.000$ $553$ $13$
R. Rosales -$0.112$ $0.000$ $556$ $13$
Goalkeeper Cuéllar -$0.048$ $0.000$ $554$ $13$
Arrizabalaga -$0.076$ $0.000$ $555$ $13$
Table 6: Selected players from each position group with regards to indirect involvement (OF) in possessions under the exclusive lasso penalty.
Score Rank
Position Player Inv OF Inv OF
Forward L. Messi $0.302$ $0.002$ $1$ $12$
J. Calleri $0.000$ -$0.409$ $9$ $556$
Midfielder Paulinho $0.000$ $0.172$ $9$ $2$
André Gomes $0.000$ $0.093$ $9$ $5$
I. Rakitić $0.000$ $0.031$ $9$ $8$
Tomás Pina $0.000$ $0.002$ $9$ $11$
Defender J. Murillo $0.000$ $0.149$ $9$ $3$
Nacho $0.000$ $0.128$ $9$ $4$
Víctor Ruíz $0.000$ $0.042$ $9$ $7$
Marcelo $0.000$ $0.023$ $9$ $9$
Luis Hernández $0.000$ -$0.008$ $9$ $552$
Goalkeeper J. Oblak $0.000$ $0.189$ $9$ $1$
Neto $0.000$ $0.050$ $9$ $6$
Fernando Pacheco $0.000$ $0.018$ $9$ $10$
Raúl Lizoáin $0.000$ -$0.009$ $9$ $553$
M. Dmitrović $0.000$ -$0.016$ $9$ $554$
Rubén $0.000$ -$0.020$ $9$ $555$
The last penalized regression approach discussed in this work is the generalized lasso. As described in Section 3.3, the shrinkage employed tries to cluster the coefficient contributions of players into groups of equal strength. The results for the penalty employed on the direct involvement coefficients are displayed in Figure 2. In order to be coherent, the players are again divided into the usual 4 position groups, although the implicit clustering of the generalized lasso does not account for position groups. The main message is consistent with the previous approaches. Offensive players perform better with regard to direct involvement than defensive players. However, most of the players are grouped into a specific sub-group of equal strength (total rank group 25 of Figure 2). Interestingly, the players selected to deviate from the average in terms of strength, i.e. the ones not grouped into the common strength group, differ from the ones selected under the exclusive lasso penalty, although not fundamentally. Furthermore, it can be noticed that the coefficients for all players are negative. This suggests that overall the contribution of average players to possessions is negative and thus also the contribution of strong players is shrunken towards a negative value. The results from the generalized lasso regularization employed on the coefficients of indirect involvement are represented in Figure 3. Once again a common strength group can be identified, which contains most of the players (total rank group 19 of Figure 3). In this case, mostly defenders and central respectively defending midfielders obtain higher scores, however, goalkeepers are primarily grouped into the cluster of average players. Finally, it has to be mentioned that the ranking is from its conceptual design to some extent similar to the exclusive lasso and thus also not ideal for ranking players in general. Nevertheless, both models provide interesting insights and the results suggest that strong (resp. weak) players can be isolated from the mass. However, the distinction between players that are average is difficult even though it may theoretically be possible within the ridge or group lasso framework.
Figure 2: Results for each of the 4 position groups with regards to direct involvement (Inv) under the generalized lasso penalty. Coefficient groups are represented by their coefficient estimate (x-axis) and their total rank (y-axis). Groups containing only one player are marked by the name of the players.
Figure 3: Results for each of the 4 position groups with regards to indirect involvement (OF) under the generalized lasso penalty. Coefficient groups are represented by their coefficient estimate (x-axis) and their total rank (y-axis). Groups containing only one player are marked by the name of the players.
4.3 Validity of ratings
As the main goal of data-driven approaches for evaluating football players is to provide an objective view of the strengths of players, it is equally necessary to objectively analyze whether a novel metric is doing a good job. This task is not straightforward as there is no ground truth for the strength of players and the opinions of practitioners and experts may vary substantially on that matter. One naive idea would be to compare the results from the novel ratings with commonly used key performance indicators such as goal scored, assists, or even more statistical metrics such as expected goals (xG, c.f. e.g. Robberechts and Davis, 2020). However, the main reason for deriving more elaborated player rating systems is to observe player strengths beyond classical performance measures, and thus a more reliable approach is desired. This section is dedicated to providing a statistical evaluation of the rating methods described in this work.
One criterion for a reasonable evaluation system of player strength discussed by Franks et al. (2016) and Hvattum and Gelade (2021) can be summarized under the term “relevance’ ’. A metric is relevant, if there is a relationship between the metric and an outcome of interest. In terms of football, the most interesting outcome is the result of a match or the number of matches won. Thus, a relevant metric should be able to establish a predictive relationship between the player strengths and the result of a match. That is, if the ratings developed under our framework are relevant, the prediction of match results should be improved by taking into account the strengths of the players involved in these matches.
In order to test the relevancy of the ratings derived in this paper, we follow an approach that is conceptually similar to the validity checks in Hvattum and Gelade (2021). However, as we only work with data from the 2017/18 season of the Spanish La Liga the framework is adjusted similarly to the one in Bajons (2023). Match outcome data of said season is divided into a training set, containing the first 280 matches of the season, and a test set, containing the remaining 100 matches. For each match, a strength estimate of each of the competing teams is given as the average over the strength values of players playing for the team. As the penalized regression models contain two different measures of strength of each player (Inv and OF), we use two distinct methods for concatenating them. First, the sum of both effects is used as discussed in Section 4.2.1. Second, the estimates are standardized such that in both cases they range from 0 (indicating the weakest player) to 1 (indicating the strongest player). Then we simply take the average of them to determine a player’s strength. Furthermore, we adapt the grouping for the exclusive lasso. As the exclusive lasso by design attempts to set most of the players in a particular group to zero, considering only the four position groups as is done in Section 4.2.1 results in too many players set exactly to zero. A more interesting grouping for predicting match results is to consider the four position groups for each specific team that is playing. In this way, the algorithm selects the strongest (and/or possibly weakest) players on each of the position groups for each team.
For prediction of match results two state-of-the-art models are considered. First, a bivariate Poisson (BP) model as discussed by Karlis and Ntzoufras (2003), which models the number of goals by the home and away team jointly based on covariates. Second, an ordered logistic regression (OLR) model as proposed in Arntzen and Hvattum (2021), which uses an ordered categorical variable with three levels as response, namely home win ($H$), draw ($D$) and away win. In particular, for both prediction models, only one covariate is used in order to analyze the respective outcomes, namely the difference in strengths of the home and away team. The strength of a team is hereby defined as the average value of the ratings as derived from each variant of our framework. The BP and OLR models are trained on the training set and the predictive performance is evaluated on the test set using proper scoring rules as suggested by Arntzen and Hvattum (2021) and Hvattum and Gelade (2021). First, the Brier score
$BS=\sum_{i=1}^{3}(d_{i}-p_{i})^{2},$ (16)
for each match in the test set is computed. Second, the informational loss is computed, which is given as
$IL=-\log_{2}\left(\sum_{i=1}^{3}p_{i}d_{i}\right),$ (17)
where $p_{i}$ is the model’s probability of outcome $i=H,D,A$ and $d_{i}=1$, if match ended with outcome $i$ and 0 otherwise. That is, for all matches in the test set, the Brier score loss and the informational loss are obtained. Thus, the evaluation of a model can be done by considering the averages of these scores over the test period. Furthermore, differences in the resulting averages can be statistically tested by using a two-tailed paired $t$-test.
Table 7: Average predictive losses from each strength model for the 100 matches in the test dataset under the different match prediction models and loss criteria.
Ordinal Logistic Bivariate Poisson
Strength Model BS IL BS IL
Ridge Sum $0.576$ $1.398$ $0.581$ $1.406$
Ridge Avg $0.574$ $1.392$ $0.579$ $\bm{1.403}$
Group Lasso Sum $\bm{0.566}$ $\bm{1.381}$ $\bm{0.578}$ $1.404$
Group Lasso Avg $0.569$ $1.387$ $0.580$ $1.410$
Excl Lasso Sum $0.577$ $1.401$ $0.589$ $1.427$
Excl Lasso Avg $0.577$ $1.403$ $0.589$ $1.427$
Gen Lasso Sum $0.611$ $1.470$ $0.618$ $1.482$
Gen Lasso Avg $0.596$ $1.443$ $0.608$ $1.465$
PCV $0.631$ $1.514$ $0.632$ $1.514$
ELO $0.638$ $1.525$ $0.614$ $1.466$
Baseline $0.636$ $1.521$ $0.637$ $1.523$
Table 7 provides the results of the predictive performance when using the respective strength estimate as a covariate. Furthermore, three reference or baseline models are included. The model termed Baseline'' is a simple model without covariate and theELO’ ’ model is derived by using ELO team ratings as taken from http://clubelo.com/. The “PCV’ ’ model uses the $PCV$ metric described in Bajons (2023). Details for the calculation of the $PCV$ value as well as a comparison of the metric with the ratings described here are found in Appendix A. The best-performing model in both prediction frameworks and under both evaluation criteria seems to be the group lasso model, however, the ridge regression model is able to keep up, even once emerging on the top. The exclusive lasso follows these two models closely and the generalized lasso performs in general worse than the aforementioned models. This result is not unexpected as we already discussed that both these models are not necessarily optimal for deriving a rating as a lot of players are merged to the average and there is no possibility to distinguish them. However, all of these models seem to perform better in terms of validity than the two reference models. Interestingly, the PCV [^4] In the initial work, Bajons (2023) uses a slightly different data setup than the one used in this paper. More details are found in Appendix A. and the ELO model both perform only slightly better than the baseline model and are outperformed by the regularized regression models. Figure 4 displays a heatmap of the p-values of individual paired $t$-tests performed on the predictive losses of the 100 matches in the test set. The results from the $t$-test indicate that at least under the OLR model the best performing models of the ridge, group lasso, and exclusive lasso penalty significantly outperform the reference models (Baseline and ELO) at the 5 % level. For the bivariate Poisson model, the differences to the baseline model are only significant at the 10 % level, but the pattern is similar. These results suggest that the regularized regression models provide a rating that from an objective viewpoint captures the important parts of the game quite well.
Figure 4: p-values of paired two-sided $t$-tests of the predictive losses of each model for 100 matches in test dataset under the different match prediction models and loss criteria.
5 Discussion
In this paper, we present a novel framework for evaluating the strength of football players. The idea of our approach is based on top-down approaches such as plus-minus (PM) models, however, we provide two novel adaptions. First, the setup relies on possession sequences, which are sequences of consecutive on-ball actions of a team. Thus, in contrast to classical PM approaches, much more granular segments of the game are used for deriving a rating, which allows for evaluating on-ball and off-ball contributions of players to the game. Second, we incorporate four different penalization structures into the regression setup. As a start, we follow the classical approach and estimate player strengths by using a ridge penalty term. This is similar to existing PM models, however analyzing possessions allows for interesting novel insights. Next, the natural grouping of players into position groups (forward, midfielder, defender, goalkeeper) is exploited via the group lasso penalty, which allows for ridge shrinkage between groups instead of all players jointly. The third penalization structure employed, the exclusive lasso, again utilizes the group structure but uses a lasso type of penalty within these, such that strong (resp. weak) players are selected, whereas average players are set to zero. Such group penalties have not been considered in the literature so far, even though it is well-known that the position of players is essential when evaluating them. As a final regularization structure, we apply the generalized lasso penalty. It implicitly tries to group players of equal strength together. This type of penalty has not yet been used for analyzing football players either and has in general not been studied very thoroughly in the context of binomial logistic regression. Therefore, we also outline a scheme for implementing such a constraint using a conic optimization approach.
The rating methods are analyzed in detail in Section 4. On the one hand, we provide comparison and football-specific implications of the methods and on the other hand, we try to objectively evaluate the validity of the rating approaches. In general, there are similarities in all the approaches. The on-ball contributions are dominated by offensive players (forwards and midfielders), while defensive players score better on the off-ball contributions. From a domain-specific viewpoint, this result is expected as the measurement for evaluating possession is whether they result in a goal or not. It is reassuring that all the methods act similarly and these results suggest that they are meaningful and informative. The distinction between active contributions and off-ball contributions can be leveraged in two ways. While it may be more reasonable to evaluate offensive players on their direct involvement contribution and defensive players on their indirect involvement, the other way around may be of interest to coaches or scouts. Considering defensive players who perform well in on-ball contributions allows for the identification of more offensively oriented defenders, which might be desirable for specific styles of play. Similarly, offensive players that demonstrate high off-ball value may have more impact on the game than their active contributions (as measured e.g. through goals, assists, etc.) are disclosing. In terms of the validity checks conducted in this study, the regularized regression models clearly and significantly outperform baseline as well as reference models. The best-performing ones are the group lasso closely followed by the ridge regression. These models are at large quite similar, however, it seems that there is a slight advantage of applying the ridge shrinkage group-wise, as is done by the group lasso. The exclusive lasso also performs fairly well in terms of validity. Moreover, it offers interesting insights that may especially benefit scouting decisions through its ability to select outstanding players of a predefined group. The generalized lasso penalty performs worst among the regularization methods, however it offers value in the interpretation of the results. The implicit grouping of players into equal-strength groups indicates that distinctions between average players have to be conducted with care.
In more general terms, while our results provide interesting and novel insights into the evaluation and the scouting of players our framework also offers the possibility to enhance plus-minus rating models via a new way of segmenting a football match as well as the usage of a variety of penalization structures. Kharrat et al. (2020) e.g. mention some pitfalls that have to be considered when employing their PM approach to evaluate players. With the possessions data at hand and grouping of players into position groups some of the pitfalls mentioned in their paper can be avoided.
Furthermore, the flexibility of our approach combined with the ease of adapting it to existing procedures yields several possible future directions to boost football player performance evaluation. For one, it might be interesting to simply adapt PM models to allow for the presented regularization schemes. Also incorporating adjustments for home advantage, manpower, league strength or relevancy of performance (cf. Kharrat et al., 2020; Pantuso and Hvattum, 2021) is straightforward. However, this becomes more relevant when a larger amount of data is available. Another interesting direction is to adapt the target variable and allow for some sort of hybrid approach. One possibility is to follow Kharrat et al. (2020) and use xG values as target variable, however, any kind of bottom-up rating that assigns a value at the end of the possession may be used. One could even think of ordinal-type target variables, however in this case the application of the more elaborated penalization structures needs to be developed further. Adapting the target may be helpful in discriminating between players, as considering goals as outcome variables leads to a sparse and imbalanced setup. Similarly, since analyzing possessions provides more detailed insights into the game than analyzing segments as defined in PM models, it is possible to adjust the outcome variable to allow for the valuation of defensive actions such as the ability of a team to stop dangerous possessions. Finally, future endeavors may explore and compare different setups for incorporating player information. In this work, we introduced the idea of taking into account on-ball as well as off-ball actions into a top-down framework. Whereas we focus on modeling the on-field contribution additively, it might be an interesting avenue to account for the off-ball involvement by separating offensive and defensive participations on the field. For instance, a player like Lionel Messi, while without a doubt positively influencing an offensive possession simply by being a threat on the field, has arguably only little effect with regards to defensive off-ball contribution.
In summary, we believe that the presented framework allows for novel perspectives and enhancements in the field of player evaluation and scouting. The results are not only from a domain-specific but also from an objective perspective useful, and thus the stakeholders involved into the sports of football such as clubs, betting companies, media as well as fans may benefit from our evaluation methods.
Appendix A Debiased machine learning models for rating players
Another approach for deriving a rating that differs from the regularized regression is presented in Bajons (2023). In this work, a debiased machine learning (DML) model for studying player strengths is introduced. The main advantage of this model over the regularized approaches is that resulting player strengths may be interpreted in a certain way, that is to some extent more intuitive than the regularized strength estimates.
A.1 The DML model and the PCV rating
Consider the following partially linear model
$Y=D\theta_{0}+g_{0}(X)+\epsilon,$ $\quad\mathbb{E}[\epsilon|D,X]=0,$ (18)
$D=m_{0}(X)+\nu,$ $\quad\mathbb{E}[\nu|X]=0,$
where $Y$ represents the outcome variable, $D$ is the treatment variable, which is potentially highly correlated with $X$, $X$ are confounders, $\theta_{0}$ is the parameter of interest, $\epsilon$ and $\nu$ are stochastic error terms. Note that $g_{0}$ and $m_{0}$ are (potentially non-linear) nuisance functions, i.e. we need to estimate them, however, this is not our primary interest. In our setup, $Y$ is the outcome variable of whether the possession resulted in a goal or not. The variable $D$ is the involvement variable of a specific player, that we want to estimate, and $X$ is the feature matrix as described in the previous section, such that $X_{mod}=(X,D)$. As $Y$ is a binary variable, the above model specification is clearly only an approximation to the true underlying data-generating process, nevertheless, it allows for nice interpretation. The estimate for $\theta_{0}$ in this case can be seen as the level shift in the probability of scoring a goal from a possession when the player corresponding to $D$ is involved in the possession.
In order to estimate the parameter $\theta_{0}$, Chernozhukov et al. (2018) propose a debiased machine learning approach, which provides means for decreasing the regularization and overfitting bias in the estimation procedure. Their approach is based on the idea of partialling out the effect of the confounders $X$ from equation (18), i.e. rewriting it in the form
$W=\nu\theta_{0}+\epsilon,$ $\quad\mathbb{E}[\epsilon|D,X]=0,$ (19)
$W=Y-\ell_{0}(X),$ $\quad\ell_{0}(X)=\mathbb{E}[Y|X]=m_{0}(X)\theta_{0}+g_{0}(X),$
$\nu=D-m_{0}(X),$ $\quad m_{0}(X)=\mathbb{E}[D|X].$
The estimation of the treatment effect $\theta_{0}$ thus amounts to executing 2 steps:
  1. 1.
    Estimate $m_{0}$ and $\ell_{0}$ using machine learning techniques of choice, i.e. predict $D$ and $Y$ using $X$.
  2. 2.
    Estimate $\theta_{0}$ by regressing the residuals $\hat{W}=Y-\hat{\ell}_{0}(X)$ onto the residual of $\hat{\nu}=D-\hat{m}_{0}(X)$.
Combining these two steps with a suitable cross-validation strategy can be shown to reduce regularization bias as well as overfitting bias, thus providing to some extent better estimates than using a direct approach for estimation of $\theta_{0}$. The above two-step procedure can be used for estimating the effect $\theta_{0}$, i.e. the procedure is repeated such that for every player $i$ in the dataset an estimate $\hat{\theta}_{0,i}$ is obtained. The resulting estimate is nicely interpretable as the level shift in the probability of scoring when player $i$ is involved. However, the raw estimate is still subject to problems that arise from the imbalance in the dataset. The effect for players with little involvement is overestimated, while the effect for players who highly participate in the game is underestimated. This is because it is more difficult to maintain a level of effectiveness when a player is a key player and receiving a lot of passes and thus exhibiting a high amount of involvements. Pure strikers for example mostly are only involved at the end of a play and do not have to make risky passes but rather convert the chances they have. Conversely, midfielders and creative players receive the ball far more often during the game. Furthermore, they have to take more risks by passing the ball through defenses or dribbling through a wall of defenders. Thus similarly estimated values of $\theta_{0}$ might hide the fact, that one player has a lot of involvements in a lot more build-up possessions than the other player and thus is much more important to the game and the team. In order to account for this fact and value players with respect to not only quality (respectively effectiveness) but also quantity, a transformation of the $\hat{\theta}_{0}$ value is proposed, which takes into account the amount of actions a player has. The metric for rating player $i$ termed possession contribution value (PCV) is given as
$PCV_{i}=N_{i}(\hat{\theta}_{0,i}+\bar{p}),$ (20)
where $N_{i}$ is the number of involvements of players $i$ in possession over the course of the season and $\bar{p}$ is the proportion of goals scored from the possessions. The rationale behind the PCV metric is the following: Considering a general possession, where no details are known about the possession, a good estimate for the probability of scoring is to simply take the proportion of scoring a goal from a possession in our dataset, so $\bar{p}$. Since $\hat{\theta}_{0,i}$ represents the level shift in the probability of scoring a goal from a possession when player $i$ is involved in the possession, then if it is known that player $i$ is involved in this general possession, adding the level shift $\hat{\theta}_{0,i}$ to $\bar{p}$ represents a more accurate estimate of the scoring probability. Multiplying this by the number of involvements results in an estimate for the average contribution to possessions by a player over the course of the observed period. Of course, a downside of such an adjustment is that players who played fewer matches during this period due to injuries, for example, might be undervalued. This effect can be alleviated by e.g. standardizing the involvements to a certain number of minutes played. Finally we note that the design matrix $X_{mod}$ in Bajons (2023) is slightly different to the one used in this paper.
A.2 Comparison of the PCV rating with the regularized regression models
Table 8 shows the top 5 players of each position group with regards to the $PCV$ value from Equation (20). As expected, forwards and midfielders obtain a better score than defenders and goalkeepers. A glance at Table 8 suggests that the results are similar to the ones from the ridge and the group lasso regression. However, Figure 5 provides a more in-depth comparison. First, it can be seen that the $PCV$ estimates are highest correlated to the direct involvement coefficients and that there seems to be very little relation between $PCV$ and the OF coefficients from the regression approaches. While this observation is in accordance with the intuition, the figure also demonstrates that there is little agreement between the regression estimates and DML ratings for defensive players. The main implication of these insights is that under the DML rating, it is more difficult to accurately evaluate the contributions of defensive players. In that sense, the regression approaches are more flexible. An advantage of the DML approach is however that the rating can be interpreted as average contribution to possessions by a player over the course of the observed period, which may be more intuitive for practitioners than the coefficients from penalized regression models.
Table 8: Top 5 of each position group (forward, midfielder, defender, goalkeeper) under the PCV rating derived from the debiased machine learning model.
Position Player PCV Score Rank
Forward L. Messi 43.23 1
L. Suárez 35.02 2
Iago Aspas 31.75 3
Cristiano Ronaldo 29.87 4
Maxi Gómez 28.30 5
Midfielder D. Wass 28.00 6
T. Kroos 27.51 7
Álex Granell 21.41 12
É. Banega 20.74 13
Samu Castillejo 20.37 14
Defender Jonny 21.61 11
C. Lenglet 15.28 38
Barragán 15.13 39
Álvaro González 14.85 41
Piqué 14.20 50
Goalkeeper M. ter Stegen 9.57 111
K. Navas 9.57 112
Adán 5.69 196
M. Dmitrović 5.07 218
J. Oblak 5.04 220
Figure 5: Correlation of the Ratings from the DML approach with the ones from ridge regression (x-axis) and the group lasso (y-axis). Each of the 4 position groups (forwards, midfielders, defenders, goalkeepers) are colored differently and each of the involvement types (Inv, OF, Sum) from the regression approaches is displayed by a different symbol.
Finally, as is observed in the main text (Section 4.3), the $PCV$ metric as derived from the DML approach performs worse than the penalization models in terms of validity. It seems that it serves mostly as a more interpretative metric, which is often preferred by practitioners. The correlation of the ratings of offensive players (forwards and midfielders) to ridge and group lasso regression indicates that at least for these groups the model provides meaningful results. In addition, the results in Bajons (2023) suggest that a different data setup may lead to improved results in terms of validity. Nevertheless, this approach is not able to match the flexibility in terms of interpretation that is gained from the regularized methods.
Appendix B Scouting players using the exclusive lasso penalty
As mentioned in the Section 3.2 the exclusive lasso offers a lot of potential for scouting players. In order to illustrate this, we consider the special case of grouping the players according to the team they play for, the possession group (as in the Section 4.2.1), and of course also the direct and indirect involvement. Since there are 20 teams playing in La Liga, this leads to $20\cdot 4\cdot 2=160$ player groups. Due to the structure of the exclusive lasso, such a grouping thus selects the most representative (i.e. the strongest resp. weakest) players for each team and on each position, all other players will obtain a coefficient of 0. For scouting this is appealing as the algorithm leaves only a range of players from each team that may be considered for recruitment. Note that teams interested in scouting players could group them in any kind of way that is of interest. Another example would be the case, where a team is interested in scouting a player playing a specific position, that is younger than a certain age and whose market value is below some budget bound. Then the team could easily form player groups accordingly while keeping other peer groups (such as team specific player position groups for example) for comparison.
Table 9: Top Players of 18 of the 20 teams of the spanish La Liga with respect to the sum of direct (Inv) and indirect (OF) involvement.
Score
Player Inv OF Sum Position Team
L. Messi 0.24604 0.15618 0.40222 Forward Barcelona
Cristiano Ronaldo 0.35760 0.00000 0.35760 Forward Real Madrid
A. Griezmann 0.26049 0.00000 0.26049 Forward Atlético Madrid
J. Murillo 0.00000 0.20968 0.20968 Defender Valencia
Juanmi 0.15745 0.03252 0.18996 Forward Real Sociedad
Iago Aspas 0.17601 0.00000 0.17601 Forward Celta de Vigo
Portu 0.17180 0.00000 0.17180 Midfielder Girona
C. Bacca 0.22732 -0.0613 0.16603 Forward Villarreal
Charles 0.15551 0.00000 0.15551 Forward Eibar
A. Sanabria 0.14040 0.00000 0.14040 Forward Real Betis
W. Ben Yedder 0.16682 -0.0401 0.12672 Forward Sevilla
Munir 0.23037 -0.10872 0.12165 Forward Deportivo Alavés
Ángel 0.11359 0.00000 0.11359 Forward Getafe
Rubén Pérez -0.03250 0.14109 0.10859 Midfielder Leganés
C. Borges 0.01045 0.08178 0.09222 Midfielder Deportivo La Coruña
E. Boateng 0.07414 0.00000 0.07414 Forward Levante
Granero 0.00000 0.06252 0.06252 Midfielder Espanyol
Iñaki Williams 0.05245 0.00000 0.05245 Forward Athletic Club
Table 9 displays the best player from each of the 20 teams of La Liga with respect to the sum of their on-ball and off-ball contributions [^5] Two teams (Málaga and Las Palmas) have only players with negative sum coefficients and thus are not depicted in this table.. From a domain-specific view, the results are reasonable and nicely interpretable. Most of the top players from the teams are forwards or midfielders. Furthermore, the best-ranked players are also the ones from the best teams that season in terms of final standings. Only from Valencia, the best player is a defender. Interestingly, Valencia played a very good season especially defensively, receiving relatively few goals (only 3 teams received fewer goals this season).
Table 10: Top 11 players selected with respect to direct involvement.
Score
Player Inv OF Sum Position Team
Cristiano Ronaldo $0.358$ $0.000$ $0.358$ Forward Real Madrid
Maxi Gómez $0.272$$0.149$ $0.124$ Forward Celta de Vigo
A. Griezmann $0.260$ $0.000$ $0.260$ Forward Atlético Madrid
L. Messi $0.246$ $0.156$ $0.402$ Forward Barcelona
Munir $0.230$$0.109$ $0.122$ Forward Deportivo Alavés
C. Bacca $0.227$$0.061$ $0.166$ Forward Villarreal
C. Stuani $0.219$$0.185$ $0.034$ Forward Girona
Lucas Pérez $0.212$$0.150$ $0.062$ Forward Deportivo La Coruña
Rodrigo $0.197$ $0.000$ $0.197$ Forward Valencia
L. Suárez $0.184$ $0.000$ $0.184$ Forward Barcelona
Iago Aspas $0.176$ $0.000$ $0.176$ Forward Celta de Vigo
When using the results for scouting or analyzing players there are however a few peculiarities to keep in mind. As described in the Section 3.2 the penalty ensures that on a group level, a ridge-like behavior is observed. That is, on the one hand, this means that in every group there is at least one representative which has a coefficient value different from zero. On the other hand, the group effect is shrunken towards zero, where more influential groups obtain less shrinkage. Within a group however, the coefficients are competitively set towards zero, such that only the most relevant group members obtain a coefficient value different from zero. For more intuition on the behavior of the exclusive lasso penalty, we refer to Campbell and Allen (2017). That being said, analyses or comparisons of individual coefficients, such as e.g. the direct involvement coefficients of specific players, have to be conducted with care. Table 10 displays the top 11 players with respect to the on-ball coefficients. Judging purely by coefficient, it seems that Cristiano Ronaldo is way stronger than his peers. However, a further look at the table shows that from two teams (Barcelona and Celta de Vigo) two forwards are in the top 11. The exclusive lasso penalty acts in a way such that the magnitude of the (shrunken) effect of a group is divided between all relevant representatives of said group. That is, if there is more than one player in a group that obtains a value different from zero, then the whole group effect is competitively shared by all of them.
Table 11: Top 5 player-team groups form the 80 direct involvement groups.
Group Effect # Group Members
Team Position Inv OF Inv OF
Celta de Vigo Forward 0.44839 -0.14855 2 1
Barcelona Forward 0.42982 0.12873 2 2
Real Madrid Forward 0.35760 0.13353 1 1
Atlético Madrid Forward 0.26049 0.10014 1 1
Deportivo Alavés Forward 0.23037 -0.07741 1 2
Table 11 shows the player-team groups with the highest magnitude of effects with respect to the direct involvement. The two best groups (the forwards from Celta de Vigo and the forwards from Barcelona) both have two selected players, whereas, for the other groups, the algorithm chose only one representative. For the two groups, the players that have coefficients different from zero compete for the magnitude of the group effect and thus it is not straightforward to compare their coefficients directly to the ones from groups with only one selected player. In order to have a fair comparison between e.g. the two selected forwards from Barcelona and the one selected forward from Real Madrid, these players would need to be merged into one group [^6] Compare e.g. with the results from Section 4.2.1..
The setting discussed in this appendix provides some intuition on the usage of the exclusive lasso. Due to the internal functionality of this penalty, there are a lot of interesting potential use cases, especially for scouting players. From the perspective of a practitioner, it is attractive that the algorithm selects only a few strong players from each group. This leads to quick identification of interesting players that may be analyzed in more detail by scouting experts and the player development staff. When trying to interpret the effects of the exclusive lasso penalty there are however pitfalls that one needs to be aware of.