### v2

# Regularized Adjusted Plus-Minus Models for Evaluating and Scouting Football (Soccer) Players using Possession Sequences

Abstract
This paper presents a novel framework for evaluating players in association football (soccer). Our method uses possession sequences, i.e.~sequences of consecutive on-ball actions, for deriving estimates for player strengths. On the surface, the methodology is similar to classical adjusted plus-minus rating models using mainly regularized regression techniques. However, by analyzing possessions, our framework is able to distinguish on-ball and off-ball contributions of players to the game. From a methodological viewpoint, the framework explores four different penalization schemes, which exploit football-specific structures such as the grouping of players into position groups as well as into common strength groups. These four models lead to four ways to rate players by considering the respective estimate of each model corresponding to the player. The ratings are used to analyze the 2017/18 season of the Spanish La Liga. We compare similarities as well as particular use cases of each of the penalized models and provide guidance for practitioners when using the individual model specifications. Finally, we conclude our analysis by providing a domain-specific statistical evaluation framework, which highlights the potential of the penalized regression approaches for evaluating players.
1 Introduction
Evaluating player performances in professional football, sometimes referred to as soccer, has attracted a lot of interest in the last decade. Due to the low-scoring and complex nature of the sport, accurately measuring a player’s contribution to the game is a challenging task. However, evaluating individual performances in team sports is a fundamental task of increasing importance. While there is an external interest from fans (McHale et al., 2012), media (Kharrat et al., 2020), and the betting industry (Michels et al., 2023) another particularly important aspect nowadays is to use player ratings to inform scouting and retention decisions of teams while being compliant with budget constraints (Payyappalli and Zhuang, 2019; Pantuso and Hvattum, 2021).
In football, the decision-making process is traditionally realized by a qualitative assessment of players by professionals such as coaches, managers, or scouts. However, lately, there have been many attempts and advances toward a more objective approach to quantifying the impact of players on the game. For example, traditional statistics such as goals or pass completion percentage, which have been used extensively to analyze the strength of players, have been replaced by metrics such as expected goals (xG, cf.e.g.Robberechts and Davis, 2020; Anzer and Bauer, 2021) or passing models based on statistical approaches (Szczepański and McHale, 2016; Håland et al., 2019, 2020). In general, these novel data-driven methods can be categorized into two distinct groups: bottom-up and top-down rating systems (cf.~Hvattum and Gelade, 2021). The former assesses player performances by assigning values to each action performed and then aggregating them for each player over the course of a relevant period (i.e.~a match or a season). Top-down ratings on the other hand evaluate players by breaking down the whole team’s performance and distributing credit onto the players involved. Typically, bottom-up approaches use very granular data on an event level in order to assign values to either specific actions such as shots (xG) or passes or more generally any type of action (cf.~Decroos et al., 2019). While bottom-up ratings are therefore able to assess a player’s quality down to each particular action performed by the athlete, they however lack the ability to account for team strengths as well as off-ball actions by teammates. In contrast, top-down ratings such as plus-minus (PM) models (see Hvattum, 2019 for an overview) rely only on information gained from specific segments of a match, which in general is much less detailed than event stream data. These approaches are by design not able to account for specific actions taken by a player. They are, however, able to incorporate effects of teammates.
In this work, we present a novel framework for evaluating player strength. On the surface, the methodology is similar to regularized adjusted plus-minus (PM) models as we mainly use penalized regression models in order to infer a rating of players. However, we extend the classical PM approach in two ways. First, we separate matches into possessions, which are sequences of consecutive on-ball actions. Thus, we consider much more granular segments of the game than existing PM models, where a segment is defined as one where the same set of players is on the field (Kharrat et al., 2020; Hvattum, 2019). The segmentation in possessions leads to a data structure where it is possible to separate effects of being part of a possession, i.e.~actually performing on-ball actions (direct involvement), from effects of only being on the pitch (indirect involvement). From a domain-specific viewpoint, this is highly attractive as it is well-known that off-ball actions have a substantial impact in soccer (cf. Spearman, 2018; Fernández et al., 2021). Using possessions data, it is also necessary to adjust the dependent variable that measures the outcome of a segment. In the classical regularized adjusted PM approach the (normalized) goal difference in a segment is used. Since possessions are way shorter segments, goal difference is not appropriate as response variable. Instead, we focus on whether the possession ends in a goal or not as our outcome variable. Second, we provide a variety of penalization structures, which have not been considered in this context. Most of the classical plus-minus models use a ridge regression penalty, i.e.~an $L_{2}$-type of penalty on the regression coefficient (cf.~e.g. Kharrat et al., 2020). In a more innovative approach, Pantuso and Hvattum (2021) adjust the penalty term by accounting for an age effect as well as a similarity effect within the shrinkage procedure. However, to the best of our knowledge, there has been no comparison or analysis of penalization structures beyond (adaptions of) the ridge penalty. We analyze several regularization structures that exploit the implicit grouping of players that arises from the different position groups on the field. In this way, our player evaluations are able to provide interesting and novel insights into player strengths where different penalties may be used for different use cases from practitioners. Finally, we conduct a statistical evaluation of our methods specifically suited for football based on validity experiments from Hvattum and Gelade (2021). Our methods are compared to reference models using different kinds of data (ELO team ratings) and another modeling approach using a debiased machine learning approach (Bajons, 2023).
The remainder of the paper is structured as follows. In Section 2 we describe the data used for this research and the necessary preprocessing steps. In Section 3, we provide details on the novel regularization structures and models used to determine player strength. Statistical descriptions are accompanied by managerial implications of each model. Section 4 presents a comparison of the ratings as obtained from the different models and our approach to statistically evaluate the methods. Finally, we conclude the paper with a discussion in Section 5.
2 Data
We use raw event stream data from the Spanish La Liga season of 2017/18, openly available via Figshare (Pappalardo et al., 2019), as a basis in order to create a dataset of possessions. A possession is defined as a sequence of consecutive on-ball events, which ends either with the opponent team gaining the possession or by an action of the referee, i.e.~a foul, the ball leaving the pitch, or the end of a period.
As the main goal is the valuation of players, it is necessary to extract information on which player was involved in a possession sequence. Extracting offensive player involvement in possessions is straightforward from the data, however, information on the opposing players is desirable as well. Since event stream data only provides details on the on-ball actions, it is difficult to extract meaningful opponent player information. One way to incorporate such knowledge is to record all players on the field during a possession. Thus, we consider a regression matrix $X_{mod}=(X_{N_{inv}},X_{N_{of}})$ with $N=N_{inv}+N_{of}$ columns and $N_{p}$ rows. Each of the involvement columns $N_{inv}$ is a binary variable for a specific player, that is 1 whenever the player was involved in the possession and 0 otherwise. Each of the $N_{of}$ columns indicates whether a specific player was on the field for the possession team (indicated by 1) or on the field for the defensive team (indicated by $-1$). In this work, $N_{inv}=N_{of}=556$ corresponds to the number of players in the data set. Note that because we are interested in evaluating players we do not extract possession-specific variables. The data includes spatial and temporal features such as the length and speed of possessions, which may help predict the success of these. However, the underlying assumption is that spatial or temporal effects are due to the players being involved in the possession as well as on the field. That is, while it may be the case that quick but long possessions increase the chance of a successful outcome it is due to the quality of players involved and on the field that it is even possible to execute such an attack.
The number of rows $N_{p}$ depends on the number of possessions analyzed. As we are interested in deriving a player rating it makes sense to only consider valuable possessions, where we define a possession as valuable if it ends in the last third of the pitch (i.e.~sufficiently close to the opponent’s goal). Following our domain-specific intuition, we further only consider possessions with at least 3 consecutive actions. The only exceptions to this rule are possessions that start with a free kick or a corner kick, as they are relevant for scoring goals and often comprise only two actions (e.g.~the corner kick and a header). This, in particular, eliminates all the penalty kicks in our data. It is common practice to exclude these highly specific situations (Robberechts and Davis, 2020; Anzer and Bauer, 2021). In most other circumstances, short possessions are a result of duels for the ball and are thus not necessarily directly relevant for detecting a player’s influence on scoring a goal from a possession. In total, the data contains $N_{p}=53233$ possessions.
Finally, a relevant response variable is extracted, which allows us to analyze the effect of players on possessions. Since goals are the most important part of the game of football, a goal indicator is taken as outcome variable. As football is a low-scoring game, one drawback is that this outcome variable is quite imbalanced. In our data, only 1.3 % of the possessions end in a goal.
On average the possessions contain $\approx 7.5$ consecutive actions with a median of $6$ consecutive actions. The average number of direct involvements for players in possession is 394.8. The distribution of the number of involvements is skewed to the right, which intuitively reflects the idea that only a few outstanding players have a lot of involvements whereas especially defenders are not as often involved in possessions. It is important to mention, that from a domain-specific viewpoint, several possible adjustments to the data setup could be considered. While we discuss some of these in more detail in Section 5, we emphasize that for the purpose of presenting the benefits of our methods, the setup is well-suited.
3 Methods and practical implications
With the possessions data at hand, the most naive and straightforward approach to deriving a player rating would be to use classic regression techniques and fit a generalized linear model to the data. Then the models’ regression coefficients could be used as a strength estimate of each player. However, as many studies point out (Gramacy et al., 2013; Sæbø and Hvattum, 2015; Hvattum, 2019; Kharrat et al., 2020), such an approach is not expected to work in the high-dimensional and sparse setup at hand, due to classical issues such as multicollinearity, overfitting and the high imbalance in outcome and independent variables.
In the following, we discuss several approaches to deriving a rating of football players from the processed event data that overcome the previously mentioned problems. The main class of models is regularized or penalized regression, often referred to as adjusted plus-minus models, which have been studied in the context of sports player ratings by many authors previously (e.g., Gramacy et al., 2013; Hvattum, 2019; Kharrat et al., 2020). We adapt the methodology to our data setup and discuss various forms of penalized regression models. For each of the models, first, a statistical description is provided and secondly, a detailed domain-specific interpretation of the penalization structure as well as implications and use cases for practitioners are discussed.
3.1 Ridge regression
One of the earliest approaches for regularization dating back to Hoerl and Kennard (1970) is ridge regression (sometimes also referred to as Tikhonov regularization). The ridge estimator is given as
[…]
where $\ell(Y;X,\beta)$ represents the (negative) log-likelihood function for the binomial response variable $Y$ in the usual generalized linear model framework that we work in. % In the case of a Gaussian likelihood function, an explicit solution for the above problem can be derived (see e.g. Hastie et al., 2013). For generalized linear models, the problem is typically formulated as a weighted least-squares problem which is then solved in an iterative fashion (cf. Friedman et al., 2010). The solution $\hat{\beta}_{ridge}$ of problem (1) is quite appealing as it shrinks the coefficients towards each other, where more shrinkage is put on the columns that explain less variance in the regression matrix $X$. That is, as we expect the response variable, i.e.~goals scored from possession, to vary most in the directions of important players, this approach shrinks regular players more than strong (or weak) players. However, players with very similar strength will likely be shrunken towards a similar value. This is quite desirable from a perspective of a rating, as every player gets a rating that reflects his strengths in relation to every other player on the field.
Another particularly nice feature arises from the data setup. As the regression matrix $X=(X_{N_{inv}},X_{N_{of}})$ consists of direct involvement columns $N_{inv}$ as well as on-field indicator columns $N_{of}$, there are two sources of information for each player $i$ represented by two estimates $\hat{\beta}_{i,inv}$ and $\hat{\beta}_{i,of}$. %The (shrunken) estimate $\hat{\beta}_{i,inv}$ of player $i$ represents a direct involvement strenght, while $\hat{\beta}_{i,of}$ characterized the effect of player $i$ being on the field. %One is the direct information of how much a player $i$ contributes to specific possessions when involved, given by the corresponding (shrunken) estimate representing player $i$’s column in the sub-matrix $X_{N_{inv}}$. The other one is the indirect contribution to possessions, given by player $i$’s estimate from the sub-matrix $X_{N_{of}}$. From a practical perspective this is nicely interpretable: A player, even though potentially not involved in the possession, usually nevertheless has an influence on the game. A threatening forward for example may create space for other players, simply because the opposing team tries to cover the player. On the other hand, a strong goalkeeper who deflects a lot of shots also clearly influences offensive possessions of the opposing team. Such effects may be measured from the estimate of indirect involvement $\hat{\beta}_{i,of}$. The total strength of a player $i$ could then e.g.~be measured by the sum of both contributions. In Section 4 we discuss several methods of assembling the two strength estimates of a player in order to obtain an overall strength estimate of players.
3.2 Group lasso and exclusive lasso
Another common form of regularization is lasso regression, where instead of imposing an $L_{2}$-norm constraint on the coefficients $\beta$ (as is done in ridge regression), an $L_{1}$-norm penalty is used. The estimate is thus defined as
[…]
This penalty allows for implicit variable selection, as some of the estimates are not only shrunken towards zero but set exactly to zero (cf.~e.g. Hastie et al., 2013). While such an effect is desirable in many situations, as variables with less influence are omitted from the regression model and only influential variables retain a value different from zero, in our setup it is suboptimal. Even though it may be possible to detect strong (or weak) players from the resulting coefficients, the majority of them will be set to zero and no interpretation is possible anymore. This is especially inconvenient for the case of recruiting players, as a team is usually looking for a tradeoff between a strong player and an affordable player. Furthermore, the appealing effect of interpreting direct and indirect involvement to possession as in the ridge regression case is not as easy to employ. In the following, we discuss variants of the lasso penalty, which are more useful for our player evaluation framework.
A particularly popular variant is the group lasso penalty (see e.g. Simon et al., 2013) given as
$\quad\min_{\beta}\ell(Y;X,\beta)+\lambda\sum_{l=1}^{m}\sqrt{p_{l}}\left\|\beta^{(l)}\right\|_{2},$ (3)
where the predictor variables are divided into $m$ groups. In equation (3), $\ell(Y;X,\beta)$ represents a scaled version of the negative log-likelihood and $p_{l}$ are weights representing the length of the $l$-th group specific estimate $\beta^{(l)}$ (cf. Simon et al., 2013). This penalty results in a shrinkage of the coefficients of complete groups to zero, i.e.~on a group level it performs shrinkage similar to lasso. However, within each group, it does not perform variable selection, leading to a similar effect as ridge regression. Thus, if each group is of size one the group lasso results in the classical lasso. On the other hand, if there is only one group the penalty is similar to ridge regression.
In our application of rating football players, there is a natural grouping of players as usually, each player plays a specific position. While these positions may vary slightly depending on specific matches and tactics used, players can be categorized into at least 4 broad categories: Forward, midfielder, defender, and goalkeeper. The group lasso enables us to distinguish between players in each of these groups. On a group level, whole groups are shrunken simultaneously. If there were one irrelevant group (say for example the goalkeepers, whose main task is not to be involved in possession of high threat), then this group would be shrunken towards zero. Within a group however, the ridge-like penalty shrinks players of similar strength to a similar value, while less shrinkage is applied to strong (or weak) players. From an interpretational viewpoint, this is highly desirable. As the tasks on the field of each individual group vary, we do not expect comparisons between two groups, e.g.~forwards and defenders, to be fair and instead might be more interested in within-group comparisons. Another dimension arises again from the distinction between direct and indirect involvement. Similar to the ridge regression case one can combine both aspects for the groups and perform within-group analyses.
Another variation of the lasso is the exclusive lasso (see e.g. Campbell and Allen, 2017), which solves the problem
$\quad\min_{\beta}\ell(Y;X,\beta)+\frac{\lambda}{2}\sum_{l=1}^{m}\left\|\beta^{(l)}\right\|_{1}^{2},$ (4)
where $\ell(Y;X,\beta)$ again represents a scaled version of an exponential family likelihood function and the regressors can be separated into $m$ different groups. The behavior of this penalty is analyzed in detail in Campbell and Allen (2017). As opposed to the group lasso, it can be obtained from equation (4) that between groups an $L_{2}$ (i.e.~ridge) penalty is applied, while within each group an $L_{1}$ (i.e.~lasso) penalty is applied. That is, on a group level the penalty does not imply shrinkage of complete groups, i.e.~in each group at least one variable will be non-zero. Within groups however only the most relevant variables will be selected, i.e.~the penalty tries to set most of the group members to zero as is usual for the lasso penalty. In the extreme case, exactly one non-zero variable of each group will be selected.
The application-specific use case of the penalty reveals more potential in scouting rather than in rating players. While one could again categorize the players into the above 4 categories based on their broad position on the pitch and analyze which players are particularly strong (resp. weak) in their respective group, the exclusive lasso motivates more interesting selection procedures. A particular use case would be a team that is interested in signing players who play a specific position and who are at a certain age level or below a certain market value. In this case, an individual group could be formed and the exclusive lasso would result in the selection of a few strong (resp. weak) players that fall into the prespecified type, while at the same time being able to account for the usual groups as defined by e.g.~player position. More details on using the exclusive lasso for scouting are found in the supplementary materials for this work.
3.3 Generalized lasso
In the previous section, we discussed penalizing structures on prespecified groups and their usage in analyzing football players. However, a different approach is to set up the penalizing structure such that implicit grouping is achieved. In the context of analyzing sports teams, Masarotto and Varin (2012) proposed the ranking lasso, given by the problem
$\min_{\beta}\ell(Y;X,\beta)+\lambda\sum_{i<j}^{N}w_{ij}|\beta_{i}-\beta_{j}|.$ (5)
The penalty term here is a variation of the lasso that allows for the grouping of the estimated abilities of players into several equivalence classes. More generally the problem can be written as
$\min_{\beta}\ell(Y;X,\beta)+\lambda||D\beta||_{1},$ (6)
where $D$ is a matrix that allows for individual specification of lasso-type penalties. The penalization structure in equation (6) is termed generalized lasso (cf.~e.g. Tibshirani and Taylor, 2011) and includes the ranking lasso as well as the more well-known fused lasso (Tibshirani et al., 2004).
Applying the generalized lasso penalty in the player evaluation setting is not completely straightforward. The aim is to cluster similar player contributions together, however as the structure of the regression matrix is $X=(X_{N_{inv}},X_{N_{of}})$, it does not make sense to consider the simple ranking lasso penalty. Instead, a more natural approach is to consider direct involvements separated from indirect ones. We first examine the structure of the matrix $D_{N_{inv}}$ for the coefficients corresponding to one of the sub-matrices of $X$, namely $X_{N_{inv}}$. The goal is to employ a penalty such that similar strength estimates are shrunken together, i.e.~to apply the ranking lasso to the estimates of these columns. Thus the matrix $D_{N_{inv}}$ can be expressed as a block matrix
$D_{N_{inv}}=(A,B,C)=\left(\begin{array}[]{ccc}A_{1}&B_{1}&C_{1}\\ \vdots&\vdots&\vdots\\ A_{N-1}&B_{N-1}&C_{N-1}\end{array}\right).$ (10)
containing three blocks $A^{\top}=(A_{1}^{\top},\dots,A_{N-1}^{\top})$, $B^{\top}=(B_{1}^{\top},\dots,B_{N-1}^{\top})$ and $C^{\top}=(C_{1}^{\top},\dots,C_{N-1}^{\top})$. Each $A_{i}$ is a matrix of zeroes of dimension $(N_{inv}-i)\times(i-1)$, each $B_{i}$ is a matrix (actually a vector) of ones of dimension $(N_{inv}-i)\times 1$ and each $C_{i}$ the negative identity matrix of dimension $N_{inv}-i$. The same penalization structure can be applied to the coefficients of indirect involvement, i.e.~these are also shrunken according to the ranking lasso penalty and thus $D_{N_{inv}}=D_{N_{of}}$. In order to obtain the full penalization matrix $D$, the two sub-matrices $D_{N_{inv}}$ and $D_{N_{of}}$ need to be concatenated suitably:
$D=\left(\begin{array}[]{cc}D_{N_{inv}}&\bm{0}\\ \bm{0}&D_{N_{of}}\end{array}\right).$ (13)
Indeed one can think of different approaches for the penalization matrix $D$, for example, if $D_{N_{of}}$ is the identity matrix, this would correspond to a classical lasso penalty on the indirect involvement coefficients. For the application to player evaluation, it makes however sense to stick to a structure where the coefficients of direct involvements in possessions are separated from the ones of indirect. In this way, this penalization approach allows for the selection of clusters of players with equal strength for on-ball and off-ball involvements separately. This may be of interest, when the goal is to identify players of similar strength and playing style compared to a reference player.
4 Computational details and results
4.1 Implementation
All of the models discussed in Section 3 are applied to the event data set described in Section 2. All computations are performed using the R statistical programming language (R Core Team, 2022) and we make use of several contributed packages. For the ridge regression, the glmnet function from the correspondent package (Friedman et al. (2010) and Tay et al., 2023) is used, the group lasso is fitted using the package SGL (Simon et al., 2019) and for the exclusive lasso the ExclusiveLasso package (Weylandt et al., 2018) is provided. In all cases, we use a suitable cross-validation procedure in order to select the optimal shrinkage parameter $\lambda$. For the generalized lasso, there is no R package readily available. Although there exists the genlasso package (Arnold and Tibshirani, 2022), this only provides an implementation for the case of Gaussian outcome variables.
We briefly describe our implementation for the binomial logistic regression. We use modern conic programming principles to solve the problem. In particular, equation (6) can be written as
$\min_{\bm{\beta}}-\big{(}\sum_{i=1}^{n}y_{i}\log(h_{\bm{\beta}}(x_{i}))+(1-y_{i})\log(1-h_{\bm{\beta}}(x_{i}))\big{)}+\lambda\sum_{j=1}^{m}w_{j}|D_{j}|,$ (14)
with logistic function $h_{\beta}(x)=1/\big{(}1+\exp(-\beta^{\top}x)\big{)}$ and pair-specific weights $w_{ij}$. In this case, $D_{j}$ is the $j$-th component of $D\bm{\beta}$, with $D\in\mathbb{R}^{m\times p}$. This problem can be reformulated as a conic program by expressing it in its equivalent epigraph form (see Boyd and Vandenberghe, 2004). Specifically, (14) can be written as
$\min_{(\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2})}$ $\sum_{i=1}^{n}t_{i}+\lambda r$ (15)
s.t. $\left(z_{i1},1,u_{i}-t_{i}\right)\in K_{\exp},$
$\left(z_{i2},1,-t_{i}\right)\in K_{\exp},$
$z_{i1}+z_{i2}\leq 1,$ $i=1,\dots,n,$
$-s_{j}\leq D_{j}\leq s_{j},$ $j=1,\dots,m,$
$r-w_{1}D_{1}-\dots-w_{m}D_{m}\geq 0.$
In the above formulation, the variables $\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2}$, and $\bm{u}$ are auxiliary variables. The initial problem is thus rewritten as a conic programming problem on an augmented set of variables $\Theta=(\bm{r},\bm{\beta},\bm{s},\bm{t},\bm{z}_{1},\bm{z}_{2})$. Such a problem can be solved using a convex optimization solver of choice. Note that in order to solve (15), only two types of conic cones are necessary, namely the linear cone and the exponential cone. We use the R-package ROI (Theußl et al., 2020) which allows to nicely handle such types of optimization problems in a flexible manner.
4.2 Comparing ratings
We provide a comparison of the penalization methods described in section 3. The aim here is to highlight idiosyncrasies as well as commonalities of each approach. A more thorough analysis containing specific rankings, player analyses, and further results can be found in the supplementary material of this work.
First, we compare the ratings obtained from ridge regression to the ones obtained from the group lasso. For the latter, the columns are split into eight groups. The players’ direct and indirect involvements are grouped into four player positions (forward, midfielder, defender, goalkeeper), resulting in four direct involvement groups and four indirect involvement groups. In general, the effect of the group lasso penalty is similar to the ridge regression as is visually confirmed by Figure 1. Offensive players (forwards and midfielders) load higher on the coefficient of direct involvement (Inv) and defensive players (defenders and goalkeepers) tend to load higher on the coefficient of indirect (on-field) involvement (OF) as is shown in the boxplots of Figure 1. However, it seems that the effects are more pronounced, especially for the direct involvements. This is reasonable since the idea of the group lasso is to shrink whole groups and thus less relevant groups, such as the defensive players, are penalized more severely. The left part of Figure 1 analyses the relationship between the ridge estimates and the group lasso estimates in more detail. In terms of correlation between the two rating methods, it seems that there is the most agreement for the coefficients of direct involvements rather than the indirect involvements. Furthermore, the ratings of the offensive players are more strongly correlated than the ones for the defensive players. Overall, the coefficients seem to be similar in both types of involvement in possessions with correlations of 0.7 or greater for all position groups. When trying to obtain a universal score for players, an intuitive approach would be to consider the sum of the coefficients from direct and indirect involvements. However, as seen in Figure 1 (left panel), the correlation between the two ratings is the lowest for the sum, suggesting that it may be worth studying different approaches of combining the two individual contributions to possession rather than simply summing them. Finally, interesting analyses can be done when comparing the Inv and OF for specific players and position groups, which we discuss in more detail in the supplement.
Figure 1: Comparison of the results from the ridge penalty and the group lasso penalty. Left: Correlation between the score of each group on the x-axis vs. correlation between the rank of each group on the y-axis. Groups are categorized by type of involvement (Inv, OF, sum) and position (FW, MF, DF, GK) Right: Boxplots of the estimates of each player in the 4 position groups under the group lasso and ridge penalty, separated by type of involvement.
Next, we discuss the results from the exclusive lasso. To allow for comparability, we use the same grouping as for the group lasso. Recall, that the exclusive lasso performs lasso-like shrinkage within each group but selects at least one relevant representative from each group. Figure 2 displays all selected players from the four direct involvement (Inv) as well as the four indirect involvement (OF) groups. The general pattern is similar to the two previous approaches. The offensive groups load stronger on Inv, with more players selected and higher coefficients (left panel of Figure 2). For OF, defensive players and groups score higher (right panel of Figure 2). As mentioned in Section 3.2, the exclusive lasso has its advantage in scouting, which would require a different setup of the groups. We refer the interested reader to the supplementary materials, where we present specific applications and discuss benefits and pitfalls when using this penalization variant.
Figure 2: Estimates of the players selected from the exclusive lasso penalty. Left: Players selected from the 4 direct involvement (Inv) position groups. Right: Players selected from the 4 indirect involvement (OF) position groups.
Finally, we analyze the results from the generalized lasso. In contrast to the group lasso and the exclusive lasso, this penalty allows for implicit grouping of players. Figure 3 shows the results for direct involvement coefficients. For comparability, we separate the players again into the usual four position groups. The main message is again consistent with the previous approaches: Offensive players perform better with regard to direct involvement than defensive players. We omit the results for indirect involvement and refer again to the supplementary material. Interestingly, the grouping formed is rather dense, i.e. outstanding players (positively and negatively) form their own groups while average players are bundled in one large group, suggesting that the distinction between these players is difficult even though theoretically possible within the ridge or group lasso framework.
Figure 3: Results for each of the 4 position groups with regards to direct involvement (Inv) under the generalized lasso penalty. Coefficient groups are represented by their coefficient estimate (x-axis) and their total rank (y-axis). Groups containing only one player are marked by the name of the players.
4.3 Validity of ratings
Having derived a statistically motivated rating, an equally important task is to objectively analyze its validity. One criterion for a reasonable evaluation system of player strength discussed by Franks et al. (2016) and Hvattum and Gelade (2021) can be summarized under the term “relevance’ ’. A metric is relevant, if there is a relationship between the metric and an outcome of interest. In terms of football, the most interesting outcome is the result of a game. Thus, a relevant metric can establish a predictive relationship between a player’s strengths and the result of a match.
The validity checks conducted in this paper follow the conceptual idea of Hvattum and Gelade (2021). %As we only work with data from the La Liga data from 2017/18, the framework is adjusted similarly to the one in Bajons (2023). In particular, match outcome data of the 2017/18 season is divided into a training set containing the first 280 matches and a test set containing the remaining 100. For each, a strength estimate of the competing teams is derived as the average over the strength values of players playing for the team. As the penalized regression models contain two different measures of player strength (Inv and OF), two distinct methods for concatenating them are applied. First, the sum of both effects is used as discussed in Section 4.2. Second, the estimates are standardized such that they range from 0 (indicating the weakest player) to 1 (indicating the strongest player) for the two contributions. Then, we average them to determine a player’s strength. Furthermore, we adapt the grouping for the exclusive lasso. As the exclusive lasso by design attempts to set most of the players in a particular group to zero, considering only the four position groups, as is done in Section~4.2, results in impractically strong sparsity. A more representative grouping for predicting match results considers the four position groups for all teams in our data. In this way, the algorithm selects the strongest (and/or possibly weakest) players on each position group for every team.
Two state-of-the-art models are considered for the prediction of match results. First, a bivariate Poisson (BP) model as discussed by Karlis and Ntzoufras (2003), which models the number of goals by the home and away team jointly based on covariates. Second, an ordered logistic regression (OLR) model as proposed in Arntzen and Hvattum (2021), which uses an ordered categorical variable with three levels as response, namely home win ($H$), draw ($D$) and away win. In particular, for both prediction models, only one covariate is used to analyze the respective outcomes, namely the difference in strengths of the home and away teams. The strength of a team is defined as the average value of the ratings as derived from each variant of our framework. Predictive performance for the models is evaluated on the test set using proper scoring rules as suggested by Arntzen and Hvattum (2021) and Hvattum and Gelade (2021). We compute the Brier score
$BS=\sum_{i=1}^{3}(d_{i}-p_{i})^{2},$ (16)
and the informational loss
$IL=-\log_{2}\left(\sum_{i=1}^{3}p_{i}d_{i}\right),$ (17)
where $p_{i}$ is the model’s probability of outcome $i=H,D,A$ and $d_{i}=1$, if match ended with outcome $i$ and 0 otherwise. That is, for all matches in the test set, the Brier score loss and the informational loss are obtained. The evaluation of a model is done by considering the averages of these scores over the test period. Furthermore, differences in the resulting averages can be statistically tested using a two-sided paired $t$-test.
Table 1: Average predictive losses from each strength model for the 100 matches in the test dataset under the different match prediction models and loss criteria.
Ordinal Logistic Bivariate Poisson
Strength Model BS IL BS IL
Ridge Sum $0.576$ $1.398$ $0.581$ $1.406$
Ridge Avg $0.574$ $1.392$ $0.579$ $\bm{1.403}$
Group Lasso Sum $\bm{0.566}$ $\bm{1.381}$ $\bm{0.578}$ $1.404$
Group Lasso Avg $0.569$ $1.387$ $0.580$ $1.410$
Excl Lasso Sum $0.577$ $1.401$ $0.589$ $1.427$
Excl Lasso Avg $0.577$ $1.403$ $0.589$ $1.427$
Gen Lasso Sum $0.611$ $1.470$ $0.618$ $1.482$
Gen Lasso Avg $0.596$ $1.443$ $0.608$ $1.465$
PCV $0.631$ $1.514$ $0.632$ $1.514$
ELO $0.638$ $1.525$ $0.614$ $1.466$
Baseline $0.636$ $1.521$ $0.637$ $1.523$
Table 1 provides the results of the predictive performance when using the respective strength estimate as a covariate. Furthermore, three reference models are included. The model termed Baseline'' is a simple model without covariate, and theELO’ ’ model is derived by using ELO team ratings as taken from http://clubelo.com/. The “PCV’ ’ model uses the $PCV$ metric described in Bajons (2023). Details for the calculation of the $PCV$ value as well as a comparison of the metric with the ratings described here are found in the supplementary material for this work. The best-performing model in both prediction frameworks and under both evaluation criteria seems to be the group lasso model. Although the ridge regression model is able to keep up, it seems that there is value in taking the positional information into account. The exclusive lasso follows these two models closely, and the generalized lasso performs in general worse than the aforementioned models. This result is not unexpected as both these models are not necessarily optimal for deriving a rating but rather have strengths in different use cases. However, all of these models perform better in terms of validity than the three reference models. Interestingly, the PCV and the ELO model both perform only slightly better than the baseline model and are outperformed by the regularized regression models. Figure 4 displays a heatmap of the p-values of individual paired $t$-tests performed on the predictive losses of the 100 matches in the test set. The results from the $t$-test indicate that at least under the OLR model the best performing models of the ridge, group lasso, and exclusive lasso penalty significantly outperform the reference models (Baseline and ELO) at the 5 % level. For the bivariate Poisson model, the differences to these models are only significant at the 10 % level, but the pattern is similar. These results suggest that the regularized regression models provide a rating that from an objective viewpoint captures the relevant parts of the game well.
Figure 4: p-values of paired two-sided $t$-tests of the predictive losses of each model for 100 matches in test dataset under the different match prediction models and loss criteria.
5 Discussion
In this paper, we present a novel framework for evaluating the strength of football players. We adapt the ideas of classical regularized adjusted plus-minus models to possession sequences data and employ a diverse set of penalization structures. The former adjustment allows for evaluating on-ball as well as off-ball contributions of players – an often overlooked but important part of the game. The latter allows for more flexibility and the incorporation of domain-specific characteristics of the game by taking into account e.g.~positions of players.
In practice, each penalization structure has its advantages respectively use cases. The general picture is, however, similar throughout all presented models. The on-ball contributions are dominated by offensive players (forwards and midfielders), while defensive players score better on the off-ball contributions. From a domain-specific viewpoint, this result is expected as the measurement for evaluating possession is whether they result in a goal or not. It is reassuring that all the methods act similarly in that sense and these results suggest that they are meaningful and informative. Furthermore, we provide objective justification for using our framework by conducting a statistically-informed tailored-to-soccer validity analysis. The results thereof suggest that using the group information of players provides value and should be accounted for. Furthermore, the additional information of on-ball and off-ball value can be leveraged in several ways. For example, defensive players who perform well in on-ball contributions may identify as offensively oriented defenders, which might be desirable for specific styles of play. Similarly, offensive players that demonstrate high off-ball value may have more impact on the game than their active contributions (as measured through goals, assists, etc.) are disclosing.
The flexibility of our approach combined with the ease of adapting it to existing procedures yields several possible future directions to boost football player performance evaluation. For one, it might be interesting to simply adapt existing regularized adjusted PM models to allow for the presented regularization schemes. Also incorporating adjustments for home advantage, manpower, league strength or relevancy of performance (cf. Kharrat et al., 2020; Pantuso and Hvattum, 2021) is straightforward. However, this becomes more relevant when more data is available. Another interesting direction is to adapt the target variable and allow for some sort of hybrid approach. One possibility is to follow Kharrat et al. (2020) and use xG values as target variable, however, any bottom-up rating that assigns a value at the end of the possession may be used. One could even think of ordinal-type target variables, however in this case the application of the more elaborated penalization structures needs to be developed further. Adapting the target may help discriminate between players, as considering goals as outcome variables leads to a sparse and imbalanced setup. Similarly, since analyzing possessions provides more detailed insights into the game than analyzing segments as defined in PM models, it is possible to adjust the outcome variable to allow for the valuation of defensive actions such as the ability of a team to stop dangerous possessions. Finally, future endeavors may explore and compare different setups for incorporating player information. For example, it might be an interesting avenue to account for the off-ball involvement by separating offensive and defensive participation on the field (indirect involvement). Figuratively speaking, a player like Lionel Messi, while without a doubt positively influencing an offensive possession simply by being a threat on the field, has arguably only little effect with regard to defensive off-ball contribution.
In summary, we believe that the presented framework allows for novel perspectives and enhancements in the field of player evaluation and scouting. The results are not only from a domain-specific but also from an objective perspective useful, and thus the stakeholders involved in the sports of football such as clubs, betting companies, media as well as fans may benefit from the presented methods.